METU Department of Mathematics Math 303 History of Math Concepts Fall 2010

Midterm 1

November 4 2010, 17:40 – 19:00

Name and the student number:

Part I

Q 1. Recognize the following numbers and determine in which civilization they were used.

Q 2. What does mean ”sexagesimal numeral system” ? Which civilization(s) was using it, and how it is used nowadays ?

Q 3. What is the name of the most famous Egyptian mathematical Papyrus ? When (approximately) it was written ? What is the name of the scribe ?

Q 4. How would Egyptian present fraction \frac{2}{5} ?

Q 5. How would Egyptians calculate the area of a circle ? Which other geometrical problems ancient Egyptians were able to solve ? (Give at least one example.)

Q 6. Write in Greek word ”theorem” and give names of letters used.

Q 7. What was the meaning of word ”mathematics” for ancient Greeks ?

Q 8. What is (approximately) time of life of Thales ? Which theorem(s) are attributed to him ?

Q 9. What is the trivium and the quadrivium ? What did they consist of ?

Q 10. What is the concept of quadrature and whom it is attributed ?

Q 11. What is the Milesian and Eleatic schools: where and by whom they were founded ?

Q 12. What are the main mathematical discoveries attributed to Pythagoras and his school ?

Q 13. Give the names of the Platonic solids and try to sketch them. What are the numbers of faces of these solids ?